Healthy people have body temperatures that are normally distributed with a mean of 98.20 degrees Fahrenheit and a standard deviation of 0.62 degrees Fahrenheit.

If a healthy person is randomly selected, what is the probability that he or she has a temperature above 98.8 degrees Fahrenheit?

A hospital wants to select a minimum temperature for requiring further medical tests. What should the temperature be, if we want only 2.5% of healthy people to exceed it?
Answer #1

μ = 98.2, σ = 0.62

(a) z = (x - μ)/σ = (98.8 - 98.2)/0.62 = 0.9677

P(x > 98.8) = P(z > 0.9677) = 0.1666

(b) z- score such that 2.5% of the area under the standard normal curve lies to its right is z = 1.96

x = μ + z * σ = 98.2 + 1.96 * 0.62 = 99.42 ° F

[Please give me a Thumbs Up if you are satisfied with my answer. If you are not, please comment on it, so I can edit the answer. Thanks.]